Given the function f(x)=√x+1 and g(x)=6x+a, in a x-y coordinate plane, y=f[g(x)] passes through the point (3,5). What is the value of 'a'?
Answer by Cromlix(4381)

f(x)=√x+1 and g(x)=6x+a
y = f(g(x)) = Square root (6x + a) + 1
If it passes thro point (3,5)
Substitute x and y values in:
y = Square root (6x + a) + 1
5 = Square root (6(3) + a) + 1
5 - 1 = Square root (18 + a)
4 = Square root (18 + a)
4^2 = 18 + a
a = 16 - 18
a = -2
a = -2
Hope this helps.
:-)